# LLMs Meet Long Video: Advancing Long Video Comprehension with An Interactive Visual Adapter in LLMs

### v1

@@ 4.5 Main Results
We present the performance of the models on four long video QA benchmarks and five short video QA benchmarks. In zero-shot long video QA benchmarks, our model achieved state-of-the-art (SOTA) results compared to the previous pure video LLMs, except WildQA. Especially on the LifeQA and Social-IQ 2.0 evaluation datasets, our model achieved significantly higher results, surpassing the previous SOTA accuracy by 18.0 and 7.4 percentage points, respectively. In zero-shot short video QA benchmarks, our model also demonstrated strong capabilities across some evaluation datasets, especially in procedure understanding. Overall, IVA significantly enhances the capability of LLMs to analyze and interpret long videos, maintaining high-performance levels without compromising the understanding and reasoning abilities of short videos.
@@ 4.6 Ablation Study
Effect of IVA. From the results in Tables 1 and 2, we found that introducing the IVA module improved the overall visual understanding of the long video datasets of Social IQ2, LifeQA, and ActivityNet-QA, as well as the short video datasets. Among them, our model achieved an improvement of over 20% on LifeQA compared to the baseline, notably suggesting the effectiveness of IVA.
@@
Figure 2: Five cases illustrate the comparative performances of our IVA Model and Baseline. The bottom part shows the detailed description of a video. Red words represent the inaccurate statement and the green words indicate the accurate statement.
Impact of long video instruction data. Furthermore, we attempted to further enhance the model’s performance during the Video Instruction Tuning phase by introducing more long video data. Therefore, we included the training sets of Social IQ2 and LifeQA, the video instruction part of the MIMIC-IT datasetLi et al. (2023a), along with the open-ended question-answering training data from NExT-QA, to the existing 170K training data, forming a new 272K training dataset. From the results of IVA(lQ=16, NI=8)-272K in the Tables 1 and 2, we observed a significant improvement on the long video dataset Social IQ2 with the inclusion of more training data. However, there was little difference in the results on the remaining datasets, and in some cases, a certain degree of decline was observed. This may be attributed to the somewhat indiscriminate addition of new datasets, leading to a certain imbalance in the proportions of different data. Additionally, the training set of LifeQA only consisted of 1,383 instances, which is relatively small in proportion to the total data, thus not providing sufficient improvement.
Number of Interaction between IVA and LLMs. Similar to the analysis of Query Tokens Length, we also conducted experiments with both doubled and halved number of interaction layers. The detailed injection layers are shown in Appendix A. Upon analyzing the results of IVA(LQ=16, NI=8) and IVA(LQ=16, NI=4) in Tables 1 and 2, we observe that this reduction resulted in a significant decrease in its performance on most long video datasets, especially on the Action Prediction and Procedure Understanding of the SEED-Bench. Moreover, the experimental results also indicate that increasing the number of layers (8 $\rightarrow$ 16) in the IVA interaction likewise caused a slight degradation in the model’s performance. Given that there was no significant improvement observed when increasing the interaction times between IVA and LLMs, we set it to 8 as the standard for experimentation.
@@ 4.7 Case Study
We present four open-ended question-answering cases and one detailed description example in Figure 2. Upon examining the initial two examples, we observe that the model augmented with IVA exhibits enhanced proficiency in recognizing particular actions associated with specific frames. In response to specific queries, it could discern objects such as the ’basketball-shaped cake’, which solely appears towards the video’s conclusion, and the ’glass bowl,’ present solely in the video’s opening segment. Furthermore, the fourth question-answering example illustrates that IVA augments the model’s reasoning ability, enabling it to deduce the prevailing weather conditions based on the lighting conditions within the video. These indicate the effectiveness of IVA in incorporating fine-grained visuals of long videos. Meanwhile, the bottom detailed description example reveals that when confronted with lengthy video descriptions, IVA could refine the perceptual acuity of LLMs, resulting in more precise recognition of elements such as the environment and color.
@@ 5 Conclusion
In this study, our primary goal is to enhance the capacity of LLMs to process and interpret long video content effectively. We identified the principal obstacles in this area and introduced an Interactive Visual Adapter (IVA) designed to facilitate dynamic interaction between LLMs and extended video sequences. The IVA incorporates a selector module for identifying relevant temporal frames within long videos based on specific instructions and tokens, along with an interactor module that isolates detailed spatial visual features within long videos. The empirical results demonstrate that our IVA significantly improves LLMs’ ability to comprehend and reason about long video content.
@@
Our work, while contributing valuable insights into video understanding through LLMs, is subject to several limitations that warrant further investigation:

### v2

@@ 4.5 Main Results
We present the performance of the models on four long video QA benchmarks and five short video QA benchmarks. In long video QA benchmarks, our model achieved state-of-the-art (SOTA) results compared to the previous pure video LLMs, except WildQA. Especially on the LifeQA and Social-IQ 2.0 evaluation datasets, our model achieved significantly higher results, surpassing the previous SOTA accuracy by 18.0 and 7.4 percentage points, respectively. In short video QA benchmarks, our model also demonstrated strong capabilities across some evaluation datasets, especially in procedure understanding. We analyze the specific performance of WildQA in Appendix B. Overall, IVA significantly enhances the capability of LLMs to analyze and interpret long videos, maintaining high-performance levels without compromising the understanding and reasoning abilities of short videos.
@@ 4.6 Ablation Study
Effect of IVA. From Tables 1 and 2, introducing the IVA module significantly improves visual understanding in long video datasets (Social IQ2, LifeQA, ActivityNet-QA) and short video datasets. Notably, our model achieved over a 20% improvement on LifeQA compared to the baseline, highlighting IVA’s effectiveness. Comparing baselines without causal or spatial tokens confirms the efficiency of our video tokens production. The experimental results in Table 3 further suggest the beneficial impact of IVA when introducing more video instruction data.
Model #DataSize SIQ2 LifeQA WildQA Avg.
Video-ChatGPT 100k 48.2/2.8 28.8/2.3 57.5/3.2 44.8
Video-LLaMA 100k 51.6/3.2 31.2/2.5 54.9/3.3 45.9
Baseline(w/o IVA) 100k 53.0/3.3 30.9/2.5 54.4/3.2 46.1
IVA (LQ=16, NI=8) 100k 53.9/3.1 33.0/2.5 57.4/3.3 48.1 $\uparrow$ 2.0
Baseline(w/o IVA) 170k 56.7/3.2 31.7/2.3 52.2/3.1 46.9
IVA (LQ=16, NI=8) 170k 64.9/3.9 50.5/3.0 53.5/3.2 56.3 $\uparrow$ 6.4
Baseline(w/o IVA) 272k 59.3/3.3 34.5/2.4 50.5/3.1 48.1
IVA (LQ=16, NI=8) 272k 68.0/4.0 48.1/2.9 50.9/3.1 55.7 $\uparrow$ 7.6
Table 3: Performance comparison when increasing the size of instruction data.
@@
[…]
Video-ChatGPT 96.83 0.9683
Video-LLaMA 127.69 1.2769
Baseline(w/o IVA) 89.18 0.8918
Baseline(IVA, LQ=16, NI=8) 90.68 0.9068
Table 4: Comparison of model efficiency. All models are tested using the same 100 samples randomly selected from the evaluation set.
Figure 2: Five cases illustrate the comparative performances of our IVA Model and Baseline. Red and green words represent the inaccurate and accurate statements, respectively.
Impact of increasing long video instruction data. We explore enhancing the model’s performance by introducing more long video data in Table 3. We use the training sets of Social IQ2 and LifeQA, the video instruction part of the MIMIC-IT datasetLi et al. (2023a), along with the open-ended QA training data from NExT-QA, to the existing 170K training data, forming a new 272K training dataset. From the results of IVA(lQ=16, NI=8)-272K in Tables 1 and 2, we observe a significant improvement in the Social IQ2 with the inclusion of more training data. However, there is little difference in the results on the remaining datasets, and in some cases, a certain degree of decline is observed. This may be attributed to the somewhat indiscriminate addition of new datasets, leading to a certain imbalance in the proportions of different data. The quality of the evaluation dataset will also affect the evaluation results, e.g., many questions in WildQA are typically wide-ranging and the criteria for answers can vary. We analyze this situation in Appendix B. Additionally, the training set of LifeQA only consists of 1,383 instances, which is relatively small in proportion to the total data, thus not providing sufficient improvement. While introducing more instruction data, IVA provides a bigger boost according to the results in Table 3.
Number of interaction between IVA and LLMs. We conduct experiments with both doubled and halved number of interaction layers. The detailed injection layers are shown in Appendix A. Upon analyzing the results of IVA(LQ=16, NI=8) and IVA(LQ=16, NI=4) in Tables 1 and 2, we observe that this reduction results in a significant decrease in its performance on most long video datasets, especially on the Action Prediction and Procedure Understanding of SEED-Bench. Moreover, the experimental results also indicate that increasing the number of layers (8 $\rightarrow$ 16) in the IVA interaction likewise causes a slight degradation in the model’s performance. Given that there is no significant improvement observed when increasing the interaction times between IVA and LLMs, we set it to 8 as the standard for experimentation.
Comparison of Computational Efficiency. We evaluate our baseline (w/o IVA), IVA (LQ=16, NI=8), and comparative models Video-ChatGPT and Video-LLaMA on 100 samples from the WildQA dataset in Table 4. All models used llama-7b or its variants with identical generation parameters. Results show: 1) Our method has the fastest inference speed, significantly surpassing Video-LLaMA. 2) With shared IVA and 8 interactions using query tokens of length 16, the increase in inference speed is marginal. Our framework outperforms previous baselines in both inference efficiency and accuracy.
@@ 4.7 Case Study
We present four Video QA cases and one detailed description example in Figure 2. Upon examining the initial two examples, we observe that the model augmented with IVA exhibits enhanced proficiency in recognizing particular actions associated with specific frames. In response to specific queries, it could discern objects such as the ’basketball-shaped cake’, which solely appears towards the video’s conclusion, and the ’glass bowl,’ present solely in the video’s opening segment. Furthermore, the fourth question-answering example illustrates that IVA augments the model’s reasoning ability, enabling it to deduce the prevailing weather conditions based on the lighting conditions within the video. These indicate the effectiveness of IVA in incorporating fine-grained visuals of long videos. Meanwhile, the bottom detailed description example reveals that when confronted with lengthy video descriptions, IVA could refine the perceptual acuity of LLMs, resulting in more precise recognition of elements such as the environment and colors.
@@ 5 Conclusion
In this study, we identify the principal obstacles in long video understanding and introduce an Interactive Visual Adapter (IVA) to facilitate dynamic interaction between LLMs and long videos. The IVA incorporates a selector module for identifying relevant temporal frames within long videos based on specific instructions and tokens, along with an interactor module that isolates detailed spatial visual features within long videos. The empirical results demonstrate that our IVA significantly improves LLMs’ ability to comprehend and reason about long video content.
@@
Our work, while contributing insights into long video understanding and question-answering through employing LLMs, is subject to several limitations that warrant further investigation:
@@ Appendix A Inserting IVA in Different Layers
Appendix B Discussion of the lower performance of IVA on the WildQA dataset.
We meticulously reviewed 652 question-and-answering pairs within the WildQA dataset, benchmarking various models against these inquiries. Of these, 319 questions were of the "what" variety, for instance, "What is the man doing?" or "What clothes is the woman wearing?" These questions are typically wide-ranging and the criteria for answers can vary, leading to inconsistencies in evaluation outcomes.
Additionally, a comparative analysis of "What"-specific accuracy versus overall sample accuracy revealed a direct correlation between IVA’s performance on "what" questions and its overall accuracy. However, IVA prioritizes visual information comprehension and its responses depend on visual elements. After introducing more data for training, as the experimental results are shown in Tables 1 and 3, improved performance across other datasets but not on WildQA further indicates the occasional disparity between its responses stemming mainly from video contents and the standard answers’ format.
Example: Question: What types of vehicles are being affected? Answer: Cars, buses, motorcycles, mopeds. Prediction: The video shows that cars and boats are being swept away by the flood water.
Model What-Acc (319 samples) All-Acc (652 samples)
Video-Chatgpt 43.88 / 2.91 54.9 / 3.3
Video-LLaMA 43.26 / 2.79 57.5 / 3.3
Baseline (w/o IVA) 40.12 / 2.74 52.2 / 3.1
IVA (LQ=16, NI=4) 40.50 / 2.80 52.5 / 3.2
IVA (LQ=16, NI=8) 41.69 / 2.78 53.5 / 3.2
IVA (LQ=16, NI=16) 42.95 / 2.83 55.1 / 3.3
IVA (LQ=16, NI=8)-272K 39.23 / 2.73 50.9 / 3.1
Table 6: Comparison of models on What-Acc and All-Acc metrics.